# Tight Robustness Certificates and Wasserstein Distributional Attacks for Deep Neural Networks

### v1

@@ 4 Wasserstein Distributional Attack
17:return $\mathbb{P}_{\mathrm{adv}}$
Figure 3: Wasserstein Distributional Attack (WDA, Alg. 1) for $r=2$. At each iteration $x_{t}$, WDA forms K1K\!-\!1 candidates $\varphi_{j}$ and updates using the one with the largest logit $\theta_{j}(\varphi_{j})$. For reference, PGD follows the dual-norm gradient direction; DeepFool linearizes the decision boundary.
5 Related Work
Robustness Certificates
Early scalable global certificates control the Lipschitz constant by multiplying per-layer operator norms, which is fast to compute yet data-agnostic and typically loose on deep nets (Virmaux & Scaman, 2018). For ReLU networks, local (activation-aware) methods exploit piecewise linearity to produce much tighter, input-conditioned certificates on individual activation regions (Katz et al., 2017; Ehlers, 2017; Weng et al., 2018; Singh et al., 2018; Shi et al., 2022). Most relevant to exact local Lipschitzness, Jordan & Dimakis (2020) showed that for a broad class of ReLU networks in general position, the local Lipschitz constant can be computed exactly by optimizing over activation patterns.
Adversarial Attacks
Adversarial Attack methods seek for perturbation $x^{\prime}$ formed by adding a small, human-imperceptible perturbation to a clean input $x$ that causes misclassification (Szegedy et al., 2014). The threat model specifies the attacker’s knowledge (white-box vs. black-box), the admissible perturbation set (e.g., $r_{2}$ balls with budget $\epsilon$), and the objective (e.g., worst-case loss within the ball). Canonical white-box methods include FGSM (Goodfellow et al., 2014), multi-step PGD (Madry et al., 2018), CW (Carlini & Wagner, 2017), and gradient-based margin attacks such as DeepFool (Moosavi-Dezfooli et al., 2016). Decision-based and score-free attacks (black-box) include Boundary Attack (Brendel et al., 2021) and Square Attack (Andriushchenko et al., 2020). Robust evaluation is subtle: gradient masking can inflate apparent robustness if attacks are not adapted (Athalye et al., 2018).
To standardize evaluation, AutoAttack (AA) (Croce & Hein, 2020) composes strong, parameter-free attacks (APGD-CE, APGD-DLR, FAB, Square) and is widely adopted for reporting robust accuracy. RobustBench (Croce et al., 2021) curates model zoos and standardized test protocols across datasets and $r_{p}$ threat models, enabling comparable and reproducible robustness claims. Liu et al. (2022) proposed Adaptive Auto Attack (A3), which incorporates Adaptive Direction Initialization (ADI) and Online Statistics-based Discarding (ODS) (Tashiro et al., 2020) to enhance attack efficiency. In our experiments, we report robustness under AA and A3 following RobustBench conventions and use them as baselines for comparison.
Several works have focused on adversarial attacks tailored to ReLU networks. Croce & Hein (2018) introduced rLR-QP, a gradient-free method that navigates the piecewise-linear regions of ReLU models by solving convex subproblems and enhancing exploration with randomization and local search. More recently, Zhang et al. (2022) developed BaB-Attack, a branch-and-bound framework that operates in activation space, leveraging bound propagation, beam search, and large neighborhood search to uncover stronger adversarial examples than conventional gradient-based approaches, particularly on hard-to-attack inputs.
6 Experiments
Experimental settings
To evaluate the effectiveness of WDA, we test the adversarial robustness of several state-of-the-art defense models on the CIFAR-10 dataset. For comparison, we report both $r_{\infty}$ and $r_{2}$ robustness under perturbation budgets of $\epsilon=8/255$ and $\epsilon=0.5$, respectively. In addition to point-wise attack, we conduct a separate Wasserstein distributional attack experiment to further validate our method. Specifically, we set $\kappa=2$ in Algorithm 1 to find the adversarial (distributional) attack $\mathbb{P}_{\mathrm{adv}}$equation 10 and reporting classification accuracy on the distribution by $(1-1/\kappa)\times\mathrm{cleanaccuracy}+\kappa\times\mathrm{adversarialaccuracy}$. Our attack is benchmarked against AA, APGD-DLR, APGD-CE (Croce & Hein, 2020), and A3 (Liu et al., 2022). The defense models, along with their official implementations and pretrained weights, are obtained from RobustBench Croce et al. (2021). All experiments are conducted on 2x NVIDIA GeForce RTX 4090 GPU and 1x NVIDIA H200.
6.1 Comparison with existing baselines
Setup
For the baselines AA, APGD-DLR, APGD-CE, and A3, we adopt the configurations reported in their respective research papers. Meanwhile, in WDA, we set the number of probe steps to $\alpha_{probe}=10$, and use $\alpha_{atk}=20$ attack iterations. We further propose A3++, an extension of A3 that incorporates our attack into its framework.
Table 1: Comparison of robust accuracy of WDA and baseline methods against various defenses on CIFAR-10 with $r_{\infty}$ perturbations ($\epsilon=8/255$) and $r_{2}$ perturbations ($\epsilon=0.5$), as well as on CIFAR-100 with $r_{\infty}$ perturbations ($\epsilon=8/255$). The first column reports the source paper corresponding to each defense method.
Paper Model Clean Single method Ensemble method
APGD -CE APGD -DLR WDA ($\kappa=1$) WDA ($\kappa=2$) AA A3 A3++
[…]
Bartoldson et al. (2024) WRN-94-16 93.68 76.15 74.31 74.05 65.25 73.71 73.55 73.54
Bartoldson et al. (2024) WRN-82-8 93.11 74.17 72.54 71.85 62.06 71.59 71.46 71.46
Cui et al. (2024) WRN-28-10 92.16 70.60 68.62 68.07 60.01 67.73 67.58 67.57
Wang et al. (2023) WRN-70-16 93.25 73.46 71.68 71.02 63.08 70.69 70.53 70.52
Wang et al. (2023) WRN-28-10 92.44 70.24 68.24 67.60 60.96 67.31 67.17 67.17
Xu et al. (2023) WRN-28-10 93.69 67.08 69.00 66.39 63.25 63.89 63.93 63.84
Sehwag et al. (2022) RN-18 84.59 58.40 57.66 56.30 54.65 55.54 55.50 55.50
[…]
Wang et al. (2023) WRN-70-16 95.54 85.66 85.30 85.00 77.63 84.97 84.96 84.97
Wang et al. (2023) WRN-28-10 95.16 84.52 83.89 83.71 76.31 83.68 83.68 83.68
Sehwag et al. (2022) WRN-34-10 90.93 78.23 78.16 77.51 72.01 77.24 77.22 77.25
Sehwag et al. (2022) RN-18 89.76 75.24 75.32 74.69 69.75 74.41 74.41 74.40
Ding et al. (2020) WRN-28-4 88.02 66.62 66.62 66.22 63.04 66.09 66.05 66.06
Cui et al. (2024) WRN-28-10 89.05 66.58 67.08 66.59 64.19 66.44 66.41 66.42
[…]
Wang et al. (2023) WRN-28-10 72.58 44.09 39.66 39.12 43.61 38.77 38.70 38.71
Addepalli et al. (2022) RN-18 65.45 33.47 28.82 28.26 37.64 27.67 27.65 27.63
Cui et al. (2024) WRN-28-10 73.85 43.82 40.37 39.57 43.68 39.18 39.17 39.14
Robustness on $r_{\infty}$ and $l_{2}$ Perturbations for CIFAR-10
Table 1 presents the robust accuracy of several attack methods under $r_{\infty}$ perturbations with $\epsilon=8/255$ and $r_{2}$ perturbations with $\epsilon=0.5$. Across both single and ensemble based threat models, WDA consistently outperforms other single-method attacks (APGD-CE and APGD-DLR), highlighting its effectiveness as a stronger standalone adversarial evaluation. Moreover, WDA achieves results that are often comparable to ensemble-based methods, indicating its ability to match the strength of more computationally demanding attack aggregations. Within the ensemble family, A3++ demonstrates clear improvements over AA and provides competitive performance with A3, surpassing it on several defense models (3 out of 7 under $r_{\infty}$ and 1 out of 6 under $r_{2}$). Notably, under $r_{\infty}$ and $r_{2}$, WDA ($\kappa=2$) produces lower robust accuracy values than other attacking methods across all defense models. This highlights the potential of Wasserstein distributional attack.
6.2 Ablation study
Figure 4: Robust accuracy with varying $\kappa$ on different defense methods.
Figure 5: Convergence of Wang et al. (2023) under $\ell_{2}$ perturbations ($\epsilon=0.5$).
Analysis on the effect of different parameter on robust accuracy
Figure 4 illustrates how the robust accuracy of WDA varies as the parameter $\kappa$ increases. Across all models, raising $\kappa$ beyond 1 generally leads to a noticeable drop in robust accuracy. Specifically, for the models from Wang et al. (2023), Sehwag et al. (2022), and Sehwag et al. (2022) (RN-18), the best performance is observed at $\kappa=3$, whereas for Ding et al. (2020), the lowest robust accuracy occurs at $\kappa=2$. These results indicate that increasing $\kappa$ consistently weakens model robustness, with the precise $\kappa$ that produces the largest drop depending on the architecture. Figure 5 presents robust accuracy for the Wang et al. (2023) model with $r_{2}$ perturbations and with different step sizes for attack. The x-axis represents iterations, and the y-axis shows robust accuracy. In Figure 5, smaller step sizes (0.01 and 0.005) lead to higher robust accuracy (97%–99%), reflecting weaker attacks. The most effective attack occurs at step size 0.2, where the accuracy drops to around 88%. At larger step sizes (0.5 and 1.0), robust accuracy stabilizes at lower values despite initial drops, suggesting reduced attack effectiveness.
@@ 7 Conclusions
We presented tight robustness certificates and stronger adversarial attacks for deep neural networks by exploiting their local geometric structure. For ReLU networks, we derived exact WDRO bounds using their piecewise-affine property, computing data-dependent Lipschitz constants from activation patterns that significantly tighten existing global bounds. For networks with smooth activations (GELU, SiLU), we characterized the worst-case loss through asymptotic Jacobian behavior along adversarial rays, providing the first tractable WDRO analysis for these modern architectures. Our Wasserstein Distributional Attack (WDA) algorithm constructs adversarial distributions on $2N$ points rather than restricting to $N$ perturbed points, achieving lower robust accuracy than state-of-the-art methods across CIFAR-10/100 benchmarks. While WDA incurs additional computational overhead compared to single-point attacks due to evaluating multiple candidate perturbations per iteration, it demonstrates that existing robustness evaluations significantly underestimate vulnerability by considering only point-wise perturbations. Together, these contributions narrow the gap between theoretical certificates and practical evaluation, revealing that both tighter bounds and stronger attacks emerge from properly leveraging network geometry and distributional perspectives.
Appendix A Proofs
A.1 Proof of Theorem 3.1
Proof of Upper Bound
@@ Appendix A Proofs
$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&=\left|\log\left[\operatorname{softmax}\theta(x^{\prime})\right]_{k}-\log\left[\operatorname{softmax}\theta(x)\right]_{k}\right|\\ &\leq 2^{1/s}\|\theta(x^{\prime})-\theta(x)\|_{s},\end{array}$ (11)
@@
$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&=\left|\left(\max_{k_{1}\neq k}\theta(x^{\prime})_{k_{1}}-\theta(x^{\prime})_{k}\right)-\left(\max_{k_{2}\neq k}\theta(x)_{k_{2}}-\theta(x)_{k}\right)\right|\\ &\leq 2^{1/s}\|\theta(x^{\prime})-\theta(x)\|_{s}.\end{array}$ (12)
In addition, by Jordan & Dimakis (2020), we have that for any $x^{\prime},x\in\mathcal{X}$,
@@
$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&\leq 2^{1/s}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}\times\|x^{\prime}-x\|_{r}\\ &=\bm{L}\times d((x^{\prime},y),(x,y)).\end{array}$ (14)
for any $x^{\prime},x\in\mathcal{X}$ and therefore by using Lipschitz certificate (Mohajerin Esfahani & Kuhn, 2018; Blanchet et al., 2019; Gao & Kleywegt, 2023; Gao et al., 2024; Chu et al., 2024), we have
@@
Proof of Lower Bound
@@
Since $\mathcal{D}_{\mathcal{X}}$ is finite, let $x^{\star},\bm{D}^{\star},k^{\prime\star},k^{\star}$ and sequence $\{u^{\star}_{t}\}$ be the maximizer in 7, i.e., $\bm{D}^{\star}\in\mathcal{D}_{x^{\star}},k^{\prime\star}\neq k^{\star},\{u^{\star}_{t}\}\subset\operatorname{int}(\operatorname{rec}(\mathcal{C}_{\bm{D}^{\star}}))$ and
@@
Choose root $\iota$ so that $(X^{(\iota)},Y^{(\iota)}=\bm{e}_{k^{\star}})$. Then when $\ell$ is the cross-entropy loss or DLR loss, by a technical Lemma A.1 one has
@@
Sufficient condition of $\bm{l}=\bm{L}$
Suppose that the dual-norm maximizer $\xi=\mathcal{M}_{r}(J_{\bm{D}^{\star}}^{\top}(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}}))\in\mathrm{rec}({\mathcal{C}_{\bm{D}^{\star}}})$ where $\bm{D}^{\star}$ is a maximizer of 6 and $(k^{\prime\star},k^{\star})$ is a maximizer of 7, then we have
@@
Figure 6: Illustration of Proof of Lower Bound
A.2 Proof of Theorem 3.3
@@
Proof of Upper Bound The WDRO upper bound is a direct consequence of the Lipschitz continuity of the loss function. The Lipschitz constant of the combined loss function, $L_{\ell}=\sup_{Z\in\mathcal{Z}}\|\nabla_{x}\ell(Z;\theta)\|_{r}$, is bounded by the product of the Lipschitz constant of the loss with respect to the output and the Lipschitz constant of the network. That is,
$\bm{L}\leq\|J(x)\|_{r\to s}\cdot\|\nabla_{\theta}\ell\|_{s}=\|J(x)\|_{r\to s}\cdot 2^{1/s}$
Proof of Lower Bound The proof of the lower bound is identical with the ReLU network case, where it relies on constructing a specific adversarial distribution. This finds a point $x^{\star}$ and a direction $u^{\star}$ that maximize the rate of change of the loss. The constant $\bm{l}$ is defined as this maximum rate of change. By constructing a perturbed point $\tilde{x}=x^{\star}+\alpha u^{\star}$ and a corresponding adversarial distribution, it is shown that the worst-case loss is at least $\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$. ∎
A.3 Technical Proofs
Lemma A.1 (Technical lemma).
In equation 19, if $\ell=\ell_{CE}$ is the cross-entropy loss, then

### v2

@@ 6 Comparison with Existing Baselines
Results on CIFAR-100 and ImageNet.
The dominance of WDA++ generalizes to larger-scale datasets. On CIFAR-100, while point-wise WDA ($\kappa=1$) remains competitive against APGD variants, WDA++ ($\Omega_{1}$) proves most effective, yielding the lowest single-method robustness. This trend continues on ImageNet, where WDA++ ($\Omega_{1}$) outperforms all baselines, demonstrating that the full distributional threat model offers the strongest evaluation of model robustness across architectures. Finally, the ensemble A3-WDA consistently sets the state-of-the-art benchmark among ensemble methods.
Runtime comparison.
Figure 4(a) benchmarks the scalability of the WDA framework against PGD baselines on ImageNet, exhibiting a linear relationship between runtime and sample size across all methods. WDA variants consistently maintain comparable or lower computational costs than their PGD counterparts.
Epsilon effect.
Figure 4(b) demonstrates that WDA++ degrades performance faster than W-PGD as $\epsilon$ increases. This stems from W-PGD’s suboptimal use of global projection, which uniformly rescales perturbations and allows high-cost samples to dilute the budget. In contrast, WDA++ employs a searching method that prioritizes cheaper samples, saving budget from hard-to-attack inputs to facilitate larger, necessary perturbations on vulnerable ones.
(a) Runtime comparison of different algorithms on the ImageNet dataset.
(b) Robust accuracy of Sehwag et al. (2022)’s ResNet-18 on CIFAR-10 under increasing $\epsilon$.
@@ 7 Conclusions
We presented tight robustness certificates and stronger adversarial attacks for deep neural networks by exploiting their local geometric structure. For ReLU networks, we derived exact WDRO bounds using their piecewise-affine property, computing data-dependent Lipschitz constants from activation patterns that significantly tighten existing global bounds. For networks with smooth activations (e.g., GELU, SiLU), we characterized the worst-case loss through asymptotic Jacobian behavior along adversarial rays, providing the first tractable WDRO analysis for these modern architectures. Our Wasserstein Distributional Attacks (WDA, WDA++) construct adversarial distributions on $2N$ points rather than restricting to $N$ perturbed points, achieving lower robust accuracy than state-of-the-art methods across CIFAR-10/100 and ImageNet benchmarks. While WDA and WDA++ incur additional computational overhead compared to single-point attacks due to evaluating multiple candidate perturbations per iteration, it demonstrates that existing robustness evaluations significantly underestimate vulnerability by considering only point-wise perturbations. Together, these contributions narrow the gap between theoretical certificates and practical evaluation, revealing that both tighter bounds and stronger attacks emerge from properly leveraging network geometry and distributional perspectives.
Appendix A Preliminaries on WDRO
Recall that given two probability measures $\mathbb{P}$ and $\mathbb{Q}$ on $\mathcal{Z}$, the Wasserstein distance is defined as
$\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\left(\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{Z}\times\mathcal{Z}}d^{p}(z^{\prime},z)\,\mathrm{d}\pi(z^{\prime},z)\right)^{1/p}$
for $p\in[1,\infty)$, and $\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\operatorname{ess.sup}_{{\pi}}(d)$ for $p=\infty$, where the feasible set is given by
$\Pi(\mathbb{P},\mathbb{Q})\triangleq\bigl\{\pi\text{ on }\mathcal{Z}\times\mathcal{Z}\mid\pi(A\times\mathcal{Z})=\mathbb{P}(A),\;\pi(\mathcal{Z}\times B)=\mathbb{Q}(B)\ \forall A,B\subseteq\mathcal{Z}\bigr\}$
the set of couplings (transport plans) between $\mathbb{P}$ and $\mathbb{Q}$. Intuitively, a transportation plan $\pi$ is feasible if it is a joint distribution whose first marginal is $\mathbb{P}$ and second marginal $\mathbb{Q}$. In the ambiguity set $\Omega_{p}=\left\{\mathbb{P}\mid\mathcal{W}_{d,p}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon\right\}$ of problem equation 3, a (distributional) attack $\mathbb{P}$ is admissible if the minimal effort for moving mass from $\mathbb{P}$ to the empirical distribution $\mathbb{P}_{N}$ is not exceeding budget $\epsilon$. Unlike traditional approaches which only allows point-wise perturbations, WDRO min-max model equation 3 allows both discrete and continuous distribution $\mathbb{P}$, which is extremely practical in certain scenarios where the ground-truth distribution $\mathbb{P}_{\rm true}$ is unknown and possibly continuous.
Appendix B Proofs of Main Results
B.1 Proof of Theorem 3.1
Proof of Upper Bound.
@@ Appendix B Proofs of Main Results
$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&=\left|\log\left(\operatorname{softmax}\theta(x^{\prime})\right)_{k}-\log\left(\operatorname{softmax}\theta(x)\right)_{k}\right|\\ &\leq 2^{1/r}\|\theta(x^{\prime})-\theta(x)\|_{s},\end{array}$ (12)
@@
$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&=\left|\left(\max_{k_{1}\neq k}\theta(x^{\prime})_{k_{1}}-\theta(x^{\prime})_{k}\right)-\left(\max_{k_{2}\neq k}\theta(x)_{k_{2}}-\theta(x)_{k}\right)\right|\\ &\leq 2^{1/r}\|\theta(x^{\prime})-\theta(x)\|_{s}.\end{array}$ (13)
In addition, by Jordan and Dimakis (2020), we have that for any $x^{\prime},x\in\mathcal{X}$,
@@
$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&\leq 2^{1/r}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}\times\|x^{\prime}-x\|_{r}\\ &=\bm{L}\times d((x^{\prime},y),(x,y)).\end{array}$ (15)
for any $x^{\prime},x\in\mathcal{X}$ and therefore by using Lipschitz certificate (Mohajerin Esfahani and Kuhn, 2018; Blanchet et al., 2019; Gao and Kleywegt, 2023; Gao et al., 2024; Chu et al., 2024), we have
@@
Proof of Lower Bound.
@@
Since $\mathcal{D}_{\mathcal{X}}$ is finite, let $x^{\star},\bm{D}^{\star},k^{\prime\star},k^{\star}$ and sequence $\{u^{\star}_{t}\}$ be the maximizer in equation 7, i.e., $\bm{D}^{\star}\in\mathcal{D}_{x^{\star}},k^{\prime\star}\neq k^{\star},\{u^{\star}_{t}\}\subset\operatorname{int}(\operatorname{rec}(\mathcal{C}_{\bm{D}^{\star}}))$ and
@@
Choose root $\iota$ so that $(X^{(\iota)},Y^{(\iota)}=\bm{e}_{k^{\star}})$. Then when $\ell$ is the cross-entropy loss or DLR loss, by a technical Lemma C.1 one has
@@
Figure 5: Illustration of Proof of Lower Bound
Sufficient condition of $\bm{l}=\bm{L}$.
Suppose that the dual-norm maximizer $\xi=\mathcal{M}_{r}(J_{\bm{D}^{\star}}^{\top}(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}}))\in\mathrm{rec}({\mathcal{C}_{\bm{D}^{\star}}})$ where $\bm{D}^{\star}$ is a maximizer of equation 6 and $(k^{\prime\star},k^{\star})$ is a maximizer of equation 7, then we have
@@
B.2 Proof of Theorem 3.3
@@
Proof of Upper Bound.
The WDRO upper bound is a direct consequence of the Lipschitz continuity of the loss function. The Lipschitz constant of the combined loss function, $L_{\ell}=\sup_{Z\in\mathcal{Z}}\|\nabla_{x}\ell(Z;\theta)\|_{r}$, is bounded by the product of the Lipschitz constant of the loss with respect to the output and the Lipschitz constant of the network. That is,
$\bm{L}\leq\|J_{\theta}(x)\|_{r\to s}\cdot\|\nabla_{\theta}\ell\|_{s}=\|J_{\theta}(x)\|_{r\to s}\cdot 2^{1/r}$
Proof of Lower Bound.
The proof of the lower bound is identical with the ReLU network case, where it relies on constructing a specific adversarial distribution. This finds a point $x^{\star}$ and a direction $u^{\star}$ that maximize the rate of change of the loss. The constant $\bm{l}$ is defined as this maximum rate of change. By constructing a perturbed point $\tilde{x}=x^{\star}+\alpha u^{\star}$ and a corresponding adversarial distribution, it is shown that the worst-case loss is at least $\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$. ∎
B.3 Proof of Corollary 3.4
Proof.
Transformer blocks are compositions of differentiable primitives (linear maps, softmax attention, residual connections, LayerNorm with $\varepsilon>0$, and smooth activations), hence differentiable on $\mathcal{X}$. To obtain equation 10, one combines explicit Jacobian bounds for softmax self-attention from prior work (Kim et al., 2021; Castin et al., 2024) with standard Lipschitz calculus: for compositions, operator norms multiply (chain rule), and for residual maps $x\mapsto x+f(x)$ one has $\mathrm{Lip}(\mathrm{Id}+f)\leq 1+\mathrm{Lip}(f)$. Applying these rules block-by-block yields a product-form upper bound for $\sup_{x\in\mathcal{X}}|J_{\theta}(x)|_{2\to 2}$, and the general $r\to s$ case follows from Hölder’s inequality. ∎
Appendix C Auxiliary Results
Lemma C.1 (Asymptotics of loss).
If $\ell=\ell_{CE}$ is the cross-entropy loss in equation 20, then